write the slope-intercept equation for the line that passes through (-1,-3) and (-2,5)
Answer by mananth(16946)

x1 y1 x2 y2
1 -3 -2 5

slope m =(y2-y1)/(x2-x1)
m=-2.67

Plug value of the slope and point (1 ,-3) in
Y=mx+b
-3.00 = -2.67 + b
b=-3-2.67
b=-0.33
So the equation will be
Y =-2.67x-0.33